Chair: Earl

Next meeting chair: Jay

Attended: Simin, Eloise, Prad, Lim, D. Schnuck, D. Smith, Jasper,

Apologies: Amin

General Issues:

Jamie has updated the slogit script to deal with the new BKG server requirements so note new syntax for slogit and flogit. Check the wiki.

Handover issues:

Hb

Two issues here:

The weather station is unresponsive. (Wth not returning valid data at the moment.)

Power-cycler for dbbc is off-line.. need to use pduhb to power-cycle. Wiki has been updated.

Ke:

Ke unstable.. but Bonn shipping new parts.

Yg:

Dbbc pps delay changed, corrupted auto-correlations…. (restart ddbc ser ver, rerun fmset, let jamie know if occurs again as could be symptomatic of more serious issue with dbbc).